# Shortcut to Chemically Accurate Quantum Computing via Density-based Basis-set Correction

### v2

@@ Abstract
Quantum computing promises a computational advantage over classical methods in electronic-structure calculations, with expected applications in drug design and materials science. Accessing a quantitative description of chemical systems while minimizing quantum resources, such as the number of qubits, is an essential challenge given the limited capabilities of current quantum processors. We provide a shortcut towards quantum computations at chemical accuracy by approaching the complete-basis-set (CBS) limit through integrating density-functional theory into quantum algorithms via density-based basis-set corrections coupled to basis-sets crafted on-the-fly and specifically adapted to a given system/user-defined qubit budget. The approach self-consistently accelerates the basis-set convergence, improving electronic densities, ground-state energies, and first-order properties such as dipole moments. It can also serve as a classical, a posteriori, energy correction to quantum hardware calculations. The strategy is assessed using GPU-accelerated state-vector emulation up to 32 qubits. We converge the ground-state energies of four atomic and molecular systems (He, Be, H2, LiH) within chemical accuracy of the CBS full-configuration-interaction reference, while offering a systematic increase of accuracy beyond a double-zeta basis-set quality for various molecules up to the H8 hydrogen chain. We also obtain dissociation curves for H2 and LiH that reach the CBS limit whereas for the challenging simulation of the N2 triple-bond breaking we achieve a near-triple-zeta basis-set quality at the cost of a minimal basis set. This hybrid strategy allows us to obtain quantitative results that would otherwise require brute-force quantum calculations using far more than 100 logical qubits, thereby opening up opportunities to explore real-world chemistry with reasonable computational resources.
@@
Quantum computing (QC) offers a promising approach to solving electronic-structure problems, with algorithms like Quantum Phase Estimation (QPE) and Variational Quantum Eigensolver (VQE) proving effective in performing ground-state quantum-chemistry wave-function calculations [1, 2, 3, 4, 5, 6]. The electronic Hamiltonian is expressed in second quantization, employing an encoding that maps one spin-orbital to one qubit. This mapping allows one to represent an exponentially large Hilbert space using only a linear number of qubits. However, to achieve accurate and practically valuable predictions for chemical systems in real-world applications, the molecular Hamiltonian should be expressed and solved using extensive basis sets of one-electron orbital functions. The number of qubits required for such calculations quickly exceeds the available capacities on current Noisy Intermediate Scale Quantum (NISQ) devices, upcoming early Fault-Tolerant Quantum Computing (FTQC) devices, and high-performance classical emulators. Therefore, while quantum chemistry has long been stated as a promising application for quantum computing, the endeavors have so far been limited to small molecular systems and minimal basis sets. However, in practice, minimal basis sets fail to be predictive for ground-state energies and chemically useful calculations require at least significantly larger double-zeta basis sets. Requirements for computing molecular properties are even more drastic as they tend to converge more slowly with the size of the basis set than energies.
@@
In electronic-structure theory, the exact solution is defined by the full-configuration interaction (FCI) method in the (infinite) complete-basis-set (CBS) limit. However, in practice, the solution derived from a finite basis set is used, which inherently suffers from truncation errors. This errors can be substantial for small basis sets, but for large enough basis sets it is possible to reach the target of chemical accuracy on energy differences, i.e. 1 kcal/mol (1.6 mHa). Unfortunately, employing sufficiently large basis sets becomes very rapidly impractical for large systems, particularly those with more than a few dozen of atoms. As a result, traditional quantum chemistry has developed a variety of computational methods to approach chemical accuracy at a reasonable cost.
@@
In particular, basis-set correction techniques such as the density-based basis-set correction (DBBSC) method [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], relying on density-functional theory (DFT), have proven effective for calculating ground- and excited-state energies, and also dipole moments, for a variety of systems including atoms, small organic molecules, and strongly correlated systems. To date, the DBBSC method has only been applied to still relatively large basis sets, more specifically the family of Dunning basis sets [20] for which chemical accuracy with respect to the CBS limit can be reached starting from a triple-zeta basis set. We note that the approximations developed within the DBBSC method target the CBS limit within a given wave-function ansatz. This means that they do not address the intrinsic errors of the wave-function method itself, such as the effects of neglected higher excitations in truncated configuration-interaction (CI) or coupled-cluster methods.
@@ I.1 Density-based basis-set correction method
In the infinite-dimensional (antisymmetric) $N$-electron Hilbert space, ${\cal H}=\bigwedge^{N}L^{2}(\mathbb{R}^{3},\mathbb{C})$, we consider an atomic or molecular system with Hamiltonian
@@ I.1.1 Strategy 1: a posteriori basis-set correction
As in previous works [7, 8], the basis-set correlation correction is evaluated at the HF density in the basis set ${\cal B}$, leading to a calculation of $\bar{E}^{\cal B}[n_{\text{HF}}^{\cal B}]$ with a marginal computational cost with respect to the cost of calculating $E_{\text{FCI}}^{\cal B}$. The PBE-based basis-set correlation correction only corrects for basis-set incompleteness errors due to short-range electron correlation. However, when using small basis sets, a significant part of the basis-set error also comes from the fact the HF part of the energy is not converged to the CBS limit. Similarly to Refs. [18, 19], we thus add a HF basis-set correction. In the present work, we choose it simply as the difference between the HF energy in the CBS limit, that we estimate as the value $E_{\text{HF}}^{\text{5Z}}$ obtained with the cc-pV5Z basis set, and the HF energy $E_{\text{HF}}^{\mathcal{B}}$ in the basis set $\mathcal{B}$
@@ I.1.2 Strategy 2: Self-consistent basis-set correction
In Strategy 2, we use the self-consistent version of the DBBSC method in Eq. (4), using the basis-set correlation correction of Eq. (5) in the self-consistent part of the calculation, and we only add a posteriori the fixed HF basis-set correction. This leads to the following approximation to the ground-state energy
@@ I.2 System-adapted basis-set generation
denote a basis set of atom-centered GTOs. It is assumed that a one-electron density, denoted by $n_{0}^{\mathcal{B}}$, is available after a converged ground-state energy minimization procedure (e.g., HF or CI). Our main purpose is to a posteriori extract subsets of the given “large” basis set $\mathcal{B}$, denoted by $\mathcal{B}_{I}=\{\chi_{\mu}\}_{\mu\in I}\subseteq\mathcal{B}$ for any index subset $I\subseteq\{1,\ldots,N_{\text{bas}}\}$, achieving (i) a target size, and (ii) minimal accuracy loss on the given density $n_{0}^{\mathcal{B}}$ used as a reference. In the present computing scheme, since the DBBSC method uses the cc-pV5Z basis set for the HF basis-set correction, we choose $\mathcal{B}$ = cc-pV5Z. In practice, this choice is motivated by the fact that such a quintuple-zeta basis set fairly approaches chemical accuracy compared to quasi-exact all-electron numeric atom-centered orbital computations [24]. Hence, given a target basis size $M$ smaller than $N_{\text{bas}}$, we seek the optimal index subset, denoted by $I_{M}$, that minimizes the best approximation error of the density $n_{0}^{\mathcal{B}}$ over the set of $\mathcal{B}_{I}$-representable one-electron densities, for any $I\subseteq\{1,\ldots,N_{\text{bas}}\}$ with $|I|=M$, where ||$\cdot$|| ⋅ | denotes the cardinal of a set, or, in other words, we solve an optimization problem under constraints:
@@
We note that the adaptation and generation of basis sets to the molecular geometry has been the subject of several publications exploring other strategies. Among them, we can cite some recent works [29, 30], but also some other research in the context of quantum computing focusing on the need to limit the qubit requirements through basis-set reoptimization such as Refs. [31, 32]. Such reoptimizations cannot be performed on the fly and obtaining such improvements can take a large amount of time when dealing with large basis sets. Alternatively, Kottman and Aspuru [33] proposed a basis-set-free approach through an adaptive representation using pair-natural orbitals which was tested up to 22 qubits.
@@ I.3 Computational Details
ADAPT-VQE computations were performed using the Hyperion-1 GPU-accelerated state-vector sparse emulator [37] up to 32 qubits. Hyperion-1 uses classical computing systems and is grounded on an efficient multi-GPU ensemble of fast custom sparse linear-algebra libraries accelerating Hyperion-1’s exact/noiseless simulations. In this paper, computations were performed on NVIDIA DGX A100 nodes (8x 80GB A100 GPUs per node) and NVIDIA DGX H100 nodes (8x 80GB A100 GPUs per node). QC calculations being strongly memory-dependent, a single GPU can carry out a 20-qubit ADAPT-VQE simulation depending on the nature (i.e. Hamiltonian sparsity) of the system whereas a single node (8 GPUs) can handle up to 28 qubits. Multi-node computations are required beyond such a qubit count. Further details about Hyperion-1 and its full capabilities will be given in a forthcoming publication. Convergence for all ADAPT-VQE computations were set to $10^{-6}$ Ha. Most computations started from a HF initial state. For selected ones (indicated in the text and Tables), we started the ADAPT-VQE procedure from a rough configuration-interaction perturbatively-selected-iteratively (CIPSI) [38] initial state (converged to only $10^{-2}$ Ha) to save some computational time within Hyperion-1. This reflects a commonly adopted strategy where a multi-determinant initial state is employed instead of a single HF determinant to increase the ground-state support in the initial state. The Quantum State Preparation of such classically-derived CIPSI wave functions has been studied in the context of VQE [39] and QPE [40]. Also, we report in Table 1, the walltime required for several ground-state energy calculations. As one can see, the walltime is not only a function of the number of qubits, sparsity is also important. For example, the Hamiltonian of the H12 molecule is sparser than the one of the water molecule resulting in a faster convergence. Beside the increased computing power of H100 GPUs leading to improved time-to-solution, our results also highlight the importance of fast node-to-node interconnects when performing large-scale quantum emulation. Indeed, the benefit of H100 over A100 is striking for the largest 32 qubits simulations on 16 nodes where an improvement of a factor 3 was observed on the DGX H100 systems. Such speedup is therefore also partially related to higher node-to-node bandwidth observed on DGX H100 versus A100 systems.
@@ II Results
We computed the ground-state energies, dissociation curves, and dipole moments for the N2, H2O, LiH, and H2 molecules using both basis-set correction strategies. For all systems except H2, the 1s molecular orbitals were frozen, i.e. we use the frozen-core approximation. Correspondingly, we use the frozen-core version of the DBBSC method as defined in Ref. [8], in which in particular the contribution to the density coming from the core electrons is neglected in the basis-set correlation density functional. The error in each calculation is quantified as the deviation from the CBS limit, which itself is determined using a two-point extrapolation scheme based on the cc-pVQZ and cc-pV5Z basis sets [21]. For the dipole-moment calculations we employed only Strategy 2, using an expectation value over the dipole-moment operator, avoiding a finite-difference approach prone to numerical errors [41].
@@ II.1 Ground-state Energies
Let us now discuss the results obtained with our SABS for LiH, N2, and H2O. They were chosen to match the sizes of the small basis sets previously discussed. Specifically for H2O, employing the V5Z-10 basis set (24 qubits) achieves results slightly superior to those obtained with the 6-31G basis set. For H2O we observed a reduction in the HF basis-set correction by 40 mHa when moving from the 6-31G to the V5Z-10 basis set, by over 100 mHa when comparing the STO-3G to the V5Z-4 basis set for the LiH molecule, and more than an Hartree for the N2 molecule moving from the STO-3G to the V5Z-6 basis set. Based on these findings, we anticipate that further exploration of this strategy could lead to a basis-set correction scheme requiring only the correlation basis-set correction term. Of course, the interest of SABS is to be systematically improvable toward the parent Dunning basis set. It is important to note that this SABS systematic improvement is not limited to quantum algorithms such as ADAPT-VQE and is also present for the HF and FCI calculations. Since the SABS approach strongly reduces the computational cost while maintaining accuracy, it should offer further applications of the CIPSI approach in classical quantum chemistry. In practice, a cc-pVTZ-like quality is reached for N2 with a V5Z-11 basis set (32 qubits) and for H2O with a V5Z-11 basis set (30 qubits). A V5Z-10 basis set (28 qubits) achieves a cc-pV5Z-like ground-state energy accuracy for LiH. Overall, the SABS always provide the best ADAPT energies with the self-consistent correlation basis-set correction (before addition of the HF basis-set correction).

### v3

@@ Abstract
Using GPU-accelerated state-vector emulation, we propose to embed a quantum computing ansatz into density-functional theory via density-based basis-set corrections (DBBSC) to obtain quantitative quantum-chemistry results on molecules that would otherwise require brute-force quantum calculations using hundreds of logical qubits. Indeed, accessing a quantitative description of chemical systems while minimizing quantum resources is an essential challenge given the limited qubit capabilities of current quantum processors. We provide a shortcut towards chemically accurate quantum computations by approaching the complete-basis-set limit through coupling the DBBSC approach, applied to any given variational ansatz, to an on-the-fly crafting of basis sets specifically adapted to a given system and user-defined qubit budget. The resulting approach self-consistently accelerates the basis-set convergence, improving electronic densities, ground-state energies, and first-order properties (e.g. dipole moments), but can also serve as a classical, a posteriori, energy correction to quantum hardware calculations with expected applications in drug design and materials science.
@@
Quantum computing (QC) offers a promising approach to solving electronic-structure problems, with algorithms like Quantum Phase Estimation (QPE) and Variational Quantum Eigensolver (VQE) proving effective in performing ground-state quantum-chemistry wave-function calculations [1, 2, 3, 4, 5, 6]. The electronic Hamiltonian is expressed in second quantization, employing an encoding that maps one spin-orbital to one qubit. This mapping allows one to represent an exponentially large Hilbert space using only a linear number of qubits. However, to achieve accurate and practically valuable predictions for chemical systems in real-world applications, the molecular Hamiltonian should be expressed and solved using extensive basis sets of one-electron orbital functions. The number of qubits required for such calculations quickly exceeds the available capacities on current Noisy Intermediate Scale Quantum (NISQ) devices, upcoming early Fault-Tolerant Quantum Computing (FTQC) devices, and high-performance classical emulators. Therefore, while quantum chemistry has long been stated as a promising application for quantum computing, the endeavors have so far been limited to small molecular systems and minimal basis sets. However, in practice, minimal basis sets fail to be predictive for ground-state energies and chemically useful calculations require at least significantly larger than double-zeta basis sets. Requirements for computing molecular properties are even more drastic as they tend to converge more slowly with the size of the basis set than energies.
@@
In electronic-structure theory, the exact solution is defined by the full-configuration interaction (FCI) method in the (infinite) complete-basis-set (CBS) limit. However, in practice, the solution derived from a finite basis set is used, which inherently suffers from truncation errors. This errors can be substantial for small basis sets, but for large enough basis sets it is possible to reach the target of chemical accuracy on energy differences, i.e. 1 kcal/mol (1.6 mHa). Unfortunately, employing sufficiently large basis sets becomes very rapidly impractical for large systems, particularly those with more than a few dozen of atoms. As a result, quantum chemists have developed a variety of traditional (i.e. classical) computational methods to approach chemical accuracy at a reasonable cost [7, 8, 9, 10]. Similarly, on the quantum computing side, the diversity of available methods has significantly increased [11, 12, 13, 14]. In particular, the hybrid quantum-classical VQE techniques initially designed to solve the general eigenvalue problem, [5] have been shown to be particularly suited for chemical applications in the present NISQ era [15, 5, 16, 15]. VQE algorithms have evolved over the years in two directions: i) fixed-length ansätze, often inspired from classical coupled cluster, with approaches such as the Unitary Coupled Cluster (UCC) and its extensions (see  [17, 18, 19] and references therein); ii) adaptive methods such as the Adaptive Derivative-Assembled Pseudo-Trotter ansatz Variational Quantum Eigensolver (ADAPT-VQE)[20] have also started to be particularly popular since they allow tailoring system-specific ansätze with shorter circuits. In recent years, ADAPT-VQE approaches have been systematically improved [21, 22, 23, 24, 25, 26] while alternative, resource-saving, adaptive techniques have also been introduced [27, 28].
However, for applications to quantum computing, the inspiration coming from classical approaches has not been limited [14] to UCC-like techniques and one can particularly mention the importance of explicitly correlated approaches [29, 30]. Indeed, several quantum strategies originate from the latter and aim to maintain full accuracy while further minimizing quantum resources for obtaining more compact wave-function representations, which is the key criterion for hardware implementations. We can particularly mention the works on transcorrelated approaches [31, 32, 33, 34, 35, 36] which introduce short-range correlation effects such as the electron-electron cusp condition [37] through various Hamiltonian modifications. Therefore, explicitly correlated quantum computing ansätze based on the transcorrelated approach have been shown to successfully evaluate ground-state energies while using fewer resources [38, 39, 40, 41, 42]. Some of them have even been extended to the evaluation of excited-state energies [39]. However, other strategies exist in classical quantum chemistry and could be particularly suited to quantum computing.
@@
In particular, basis-set correction techniques such as the density-based basis-set correction (DBBSC) method [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55], relying on density-functional theory (DFT), have proven to be effective for calculating ground- and excited-state energies, and also dipole moments, for a variety of systems including atoms, small organic molecules, and strongly correlated systems. This approach provides the key benefit of an accelerated convergence to the CBS limit with the basis-set size, which is a valuable asset for quantum computing since minimizing quantum resources is paramount. To date, the DBBSC method has only been applied to still relatively large basis sets, more specifically the family of Dunning basis sets [56] for which chemical accuracy with respect to the CBS limit can be reached starting from a triple-zeta basis set. We note that the approximations developed within the DBBSC method target the CBS limit within a given wave-function ansatz. This means that they do not address the intrinsic errors of the wave-function method itself, such as the effects of neglected higher-excitations in truncated configuration-interaction (CI) or coupled-cluster methods.
@@ I.1 Density-based basis-set correction method
In the infinite-dimensional (antisymmetric) $N$-electron Hilbert space, ${\cal H}=\bigwedge^{N}L^{2}(\mathbb{R}^{3}\times\{\uparrow,\downarrow\},\mathbb{C})$, we consider an atomic or molecular system with Hamiltonian
@@ I.1.1 Strategy 1: a posteriori basis-set correction
As in previous works [43, 44], the basis-set correlation correction is evaluated at the HF density in the basis set ${\cal B}$, leading to a calculation of $\bar{E}^{\cal B}[n_{\text{HF}}^{\cal B}]$ with a marginal computational cost with respect to the cost of calculating $E_{\text{FCI}}^{\cal B}$. The PBE-based basis-set correlation correction only corrects for basis-set incompleteness errors due to short-range electron correlation. However, when using small basis sets, a significant part of the basis-set error also comes from the fact the HF part of the energy is not converged to the CBS limit. Similarly to Refs. [54, 55], we thus add a HF basis-set correction. In the present work, we choose it simply as the difference between the HF energy in the CBS limit, that we estimate as the value $E_{\text{HF}}^{\text{5Z}}$ obtained with the cc-pV5Z basis set [59], and the HF energy $E_{\text{HF}}^{\mathcal{B}}$ in the basis set $\mathcal{B}$
@@
Let us emphasize again that the PBE-based basis-set correction mentioned above only takes care of the correlation part and therefore does not correct the basis-set error of the HF energy. Hence, there is no double counting between the two basis-set corrections, as shown in Refs. [55, 54].
@@ I.1.2 Strategy 2: Self-consistent basis-set correction
In Strategy 2, we use the self-consistent version of the DBBSC method in Eq. (4), using the basis-set correlation correction of Eq. (5) in the self-consistent part of the calculation, and we only add a posteriori the fixed HF basis-set correction. This leads to the following approximation to the ground-state energy that can be applied to any variational quantum ansatz
@@
In Strategy 2, the density of the QC wave-function ansatz $\Psi$ needs to be calculated. For this, we calculate the one-particle density matrix
$n_{pq,\sigma}=\langle\Psi,\hat{a}^{\dagger}_{p,\sigma}\hat{a}_{q,\sigma}\Psi\rangle,$ (13)
where $\hat{a}^{\dagger}_{p,\sigma}$ and $\hat{a}_{q,\sigma}$ are the creation and annihilation operators, respectively, for the $p$th and $q$th orbitals and with spin $\sigma\in\{\uparrow,\downarrow\}$. This is achieved with a Jordan-Wigner mapping of the operator $\hat{a}^{\dagger}_{p,\sigma}\hat{a}_{q,\sigma}$ and the state $\Psi$ prepared using the parameterized ansatz. The density matrix is then used to classically update the Hamiltonian in Eq.(10) and to calculate dipole moments.
@@ I.2 System-adapted basis-set generation
denote a spatial basis set of atom-centered GTOs. It is assumed that a one-electron density, denoted by $n_{0}^{\mathcal{B}}$, is available after a converged ground-state energy minimization procedure (e.g., HF or CI). Our main purpose is to a posteriori extract subsets of the given “large” basis set $\mathcal{B}$, denoted by $\mathcal{B}_{I}=\{\chi_{\mu}\}_{\mu\in I}\subseteq\mathcal{B}$ for any index subset $I\subseteq\{1,\ldots,N_{\text{bas}}\}$, achieving (i) a target size, and (ii) minimal accuracy loss on the given density $n_{0}^{\mathcal{B}}$ used as a reference. In the present computing scheme, since the DBBSC method uses the cc-pV5Z basis set for the HF basis-set correction, we choose $\mathcal{B}$ = cc-pV5Z. In practice, this choice is motivated by the fact that such a quintuple-zeta basis set fairly approaches chemical accuracy compared to quasi-exact all-electron numeric atom-centered orbital computations [61]. Hence, given a target basis size $M$ smaller than $N_{\text{bas}}$, we seek the optimal index subset, denoted by $I_{M}$, that minimizes the best approximation error of the density $n_{0}^{\mathcal{B}}$ over the set of $\mathcal{B}_{I}$-representable one-electron densities, for any $I\subseteq\{1,\ldots,N_{\text{bas}}\}$ with $|I|=M$, where ||$\cdot$|| ⋅ | denotes the cardinal of a set, or, in other words, we solve an optimization problem under constraints:
@@
We note that the adaptation and generation of basis sets to the molecular geometry has been the subject of several publications exploring other strategies. Among them, we can cite some recent works [66, 67], but also some other research in the context of quantum computing focusing on the need to limit the qubit requirements through basis-set reoptimization such as Refs. [68, 69]. Alternatively, Kottman and Aspuru [70] proposed a basis-set-free approach through an adaptive representation using pair-natural orbitals which was tested up to 22 qubits.
@@ I.3 Computational Details
ADAPT-VQE computations were performed using the Hyperion-1 GPU-accelerated state-vector sparse emulator [73] up to 32 qubits. Hyperion-1 uses classical computing systems and is grounded on an efficient multi-GPU ensemble of fast custom sparse linear-algebra libraries accelerating Hyperion-1’s exact/noiseless simulations. In this paper, computations were performed on NVIDIA DGX A100 nodes (8x 80GB A100 GPUs per node) and NVIDIA DGX H100 nodes (4x & 8x 80GB H100 GPUs per node). QC calculations being strongly memory-dependent, a single GPU can carry out a 20-qubit ADAPT-VQE simulation depending on the nature (i.e. Hamiltonian sparsity) of the system whereas a single node (8 GPUs) can handle up to 28 qubits. Multi-node computations are required beyond such a qubit count. Further details about Hyperion-1 and its full capabilities will be given in a forthcoming publication. Convergence for all ADAPT-VQE computations were set to $10^{-6}$ Ha. Most computations started from a HF initial state. For selected ones (indicated in the text and Tables), we started the ADAPT-VQE procedure from a rough configuration-interaction perturbatively-selected-iteratively (CIPSI) [74] initial state (converged to only $10^{-2}$ Ha) to save some computational time within Hyperion-1. This reflects a commonly adopted strategy where a multi-determinant initial state is employed instead of a single HF determinant to increase the ground-state support in the initial state. The Quantum State Preparation of such classically-derived CIPSI wave functions has been studied in the context of VQE [25] and QPE [26]. Also, we report in Table 1, the walltime required for several ground-state energy calculations. As one can see, the walltime is not only a function of the number of qubits, sparsity is also important. For example, the Hamiltonian of the H12 molecule is sparser than the one of the water molecule resulting in a faster convergence. Beside the increased computing power of H100 GPUs leading to improved time-to-solution, our results also highlight the importance of fast node-to-node interconnects when performing large-scale quantum emulation. Indeed, the benefit of H100 over A100 is striking for the largest 32 qubits simulations on 16 nodes where an improvement of a factor 3 was observed on the DGX H100 systems. Such speedup is therefore also partially related to higher node-to-node bandwidth observed on DGX H100 versus A100 systems.
@@ II Results
We computed the ground-state energies, dissociation curves, and dipole moments for the N2, H2O, LiH, and H2 molecules using both basis-set correction strategies. For all systems except H2, the 1s molecular orbitals were frozen, i.e. we use the frozen-core approximation. Correspondingly, we use the frozen-core version of the DBBSC method as defined in Ref. [44], in which in particular the contribution to the density coming from the core electrons is neglected in the basis-set correlation density functional. The error in each calculation is quantified as the deviation from the CBS limit, which itself is determined using a two-point extrapolation scheme based on the cc-pVQZ and cc-pV5Z basis sets [57]. For the dipole-moment calculations we employed only Strategy 2, using an expectation value over the dipole-moment operator, avoiding a finite-difference approach prone to numerical errors [75]. We remind the reader that these calculations are to be interpreted in the framework of perfect (noiseless) logical qubits, and therefore the discussed errors can be only attributed to the method used. Table 4 in Appendix presents the quantum circuits that were used for all computations and provides the number of CNOT gates required to perform for the algorithm as well as the number of ADAPT-VQE iterations.
@@ II.1 Ground-state Energies
Let us now discuss the results obtained with our SABS for LiH, N2, and H2O. They were chosen to match the sizes of the small basis sets previously discussed. Specifically for H2O, employing the V5Z-10 basis set (24 qubits) achieves results slightly superior to those obtained with the 6-31G basis set. For H2O we observed a reduction in the HF basis-set correction by 40 mHa when moving from the 6-31G to the V5Z-10 basis set, by over 100 mHa when comparing the STO-3G to the V5Z-4 basis set for the LiH molecule, and more than an Hartree for the N2 molecule moving from the STO-3G to the V5Z-6 and V5Z-11 basis sets. Based on these findings, we anticipate that further exploration of this strategy could lead to a basis-set correction scheme requiring only the correlation basis-set correction term. Of course, the interest of SABS is to be systematically improvable toward the parent Dunning basis set. It is important to note that this SABS systematic improvement is not limited to quantum algorithms such as ADAPT-VQE and is also present for the HF and FCI calculations. Since the SABS approach strongly reduces the computational cost while maintaining accuracy, it should offer further applications of the CIPSI approach in classical quantum chemistry. In practice, a cc-pVTZ-like quality is reached with a V5Z-11 basis set (32 qubits) for N2, and with a V5Z-11 basis set (30 qubits) for H2O. A V5Z-10 basis set (28 qubits) achieves a cc-pV5Z-like ground-state energy accuracy for LiH. Overall, the SABS always provide the best ADAPT energies with the self-consistent correlation basis-set correction (before addition of the HF basis-set correction).